OpenStudy (anonymous):

A food store makes a 11–pound mixture of peanuts, almonds, and raisins. The cost of peanuts is $1.50 per pound, almonds cost $3.00 per pound, and raisins cost $1.50 per pound. The mixture calls for twice as many peanuts as almonds. The total cost of the mixture is $21.00. How much of each ingredient did the store use 3lbs peanuts, 6 lbs almonds, 2 lbs raisins 8 lbs peanuts, 1 lb almonds, 2 lbs raisins 6 lbs peanuts, 3 lbs almonds, 2 lbs raisins 8 lbs peanuts, 2 lbs almonds, 1 lbs raisins

OpenStudy (aripotta):

c

OpenStudy (anonymous):

Can you explain how you got it? please

OpenStudy (aripotta):

i can try, it'll take some time to type up though :/

OpenStudy (anonymous):

well its okay, you helped enough (:

OpenStudy (aripotta):

system of equations. they want 2 times the amount of almonds for peanuts. 1.5(2a) + 3a + 1.5r = 21 2a + a + r = 11 solve for r in the second equation: 3a + r = 11 r = 11 - 3a substitute for r in the first equation: 1.5(2a) + 3a + 1.5(11 - 3a) = 21 simplify/distributive property: 3a + 3a + 16.5 - 4.5a = 21 combine like terms: 1.5a + 16.5 = 21 subtract 16.5 from both sides: 1.5a = 4.5 divide by 1.5 on both sides: a = 3 and the only choice with 3 for almonds is c.

OpenStudy (aripotta):

does that make sense? and sorry it took so long

OpenStudy (aripotta):

@alyssacherry2 ._.

OpenStudy (anonymous):

sorry cooking chicken lmao! @AriPotta

OpenStudy (aripotta):

oh ok, sorry lol

OpenStudy (anonymous):

No problem (: thank you sooo much for explaining it (:

OpenStudy (aripotta):

does it make sense or do you need me to explain anything better?

OpenStudy (anonymous):

No i get it (: i can come back to this and look what you put as help fo new problems (: